OpenStudy (anonymous):

1. Kevin ran 4 miles more than Steve ran. The sum of their distances is 26 miles. How far did Steve run? The domain of the solution is {0, 4, 11, 22}. a. Define your variable. b. Write algebraic expressions for the distances each person ran. c. Write an equation that can help you solve the problem. d. Solve the equation. Check your answer

OpenStudy (anonymous):

X= Kevin and Y= Steve

OpenStudy (anonymous):

x+y=26 Because 26 is the total amount of miles

OpenStudy (anonymous):

x=4+y Kevin= 4+ Steve

OpenStudy (anonymous):

We know how many more miles Kevin ran than Steve, now we need to find out the amount of miles Steve ran so we can figure out how many miles Kevin ran

OpenStudy (anonymous):

Sub for Kevin (x) (y+4)+y=26

OpenStudy (anonymous):

2y+4=26

OpenStudy (anonymous):

Now subtract 4 from each side

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

that'll give yhu 2y=22

OpenStudy (anonymous):

but i already got the answer

OpenStudy (anonymous):

ok